OpenStudy (anonymous):

Limit of (sin(x-5))/((X^2)-25) when x approaches 5.

OpenStudy (anonymous):

\[\frac{\sin(x-5)}{(x+5)(x-5)}=\frac{\sin(x-5)}{x-5}\times \frac{1}{x+5}\] is a start

OpenStudy (anonymous):

Ok so then I get lim x -> 5 ((sin(x-5))/(x-5)) * lim x -> 5 (1/(x+5)) then lim x -> 5 ((sin(x-5))/(x-5)) * (1/10)